OpenStudy (anonymous):

Find the equation of a sine function that has a vertical displacement 2 units down, a horizontal phase shift 60° to the right, a period of 30°, a reflection in the y-axis and an amplitude of 3.

OpenStudy (anonymous):

@callisto can you help here please?

OpenStudy (anonymous):

Im working on it but Im not sure if its going the right path. Could the end of the new equation be 2(x+30degrees)-2?

OpenStudy (anonymous):

Could it be 3 sin (30(x)- 60) -2 ?

OpenStudy (callisto):

sine function -> y=sinx vertical displacement 2 units down, => y = sinx -2 a horizontal phase shift 60° to the right -> y = sin(x-60) -2 period of 30° = 1/12 (360) -> y = sin(12x-60) -2 a reflection in the y-axis -> y = sin(-12x-60) -2 = -sin(12x+60)-2 an amplitude of 3 -> y=-3sin(12x+60)-2 Hmmm... I'm really not good at graphs :(

OpenStudy (anonymous):

Thank you. I will just examine the graph:)

OpenStudy (callisto):

Hmm.. perhaps wolf helps? not sure though

OpenStudy (anonymous):

Ok, so your graph does have the correct properties

OpenStudy (anonymous):

The amplitude is 3

OpenStudy (anonymous):

I understand most of the equation, but can you explain the period of 30° = 1/12 (360) -> y = sin(12x-60) -2 part?

OpenStudy (callisto):

Hmm.. since the period is 30, which is reduced by 1/12 of the original sine graph. Like the previous one, if you reduce by 1/2, you would write y=sin (2x). Now, it's 12, so, multiply x by 12. I think it would work... Not sure though :S

OpenStudy (anonymous):

Hey, In the geogebra site, is there a way I can see degrees in the grid instead of numbers?